OpenStudy (anonymous):

Tickets at a play cost $10 or $14. If 800 tickets were sold and $9,040 was collected by the box office, how many $14 tickets were sold?

OpenStudy (anonymous):

These answers are possible! (A) 100 (B) 260 (C) 540 (D) 710

OpenStudy (anonymous):

You just have to find the right one! @TestTakingLifesaver

OpenStudy (anonymous):

Let x be $10 tickets and let y be $14 tickets. x+y=800 10x+14y=9,040 Now you have a system of equations to solve

OpenStudy (anonymous):

100??

OpenStudy (anonymous):

Nope, do you know how to solve systems of equations or do you want me to get you started

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

x+y=800 10x+14y=9,040 ----------------- we wanna multiply one of the equations by a number so that when we add them together, the x will cancel out and we can solve for y. So lets multiply the first one by -10. -10x+-10y=-8000 10x+14y=9040 -----------------add them together 4y=1040 solve for y